# Global decomposition of networks into multiple cores formed by local hubs

### v1

@@ III.2 Core-periphery structure of communities and the supernode network
Figure 7: (a) An example of a network with a community structure marked by different background colors. (b) Inside an arbitrary community of the CB network, we plot the connection densities of core-core, core-periphery, and periphery-periphery, along with the core-periphery score, as a function of the boundary level $L_{B}$. The same colors and shapes are used for the densities and the score as in Fig. 6. The core-periphery score $S_{cp}$ reaches its maximum at $L_{B}=2$. (c) The adjacency matrix represents the connections (yellow in the black background) inside the community in panel (b), where nodes are sorted by their LED levels (higher levels toward the left bottom). The nodes inside the red square are the optimal core nodes of this community.
@@
Figure 8: (a) The supernode network is composed of the nodes corresponding to the communities in the network shown in Fig. 7(a). (b) The connection densities of core-core, core-periphery, and periphery-periphery, along with the core-periphery score, as a function of the boundary level $L_{B}$ in the supernode network from the CB network. The same colors and shapes are used for the densities and the score as in Figs. 6 and 7. The core-periphery score $S_{cp}$ reaches its maximum at $L_{B}=4$. (c) The adjacency matrix represents the connections (green in the black background) between the supernodes in panel (b), where nodes are sorted by their LED levels (higher levels toward the left bottom). The supernodes inside the red square are the optimal core nodes of this community.
@@
In the previous subsection, we have found a single core of the network, but as introduced in the first part of Sec. III, more recent studies on the CP emphasize the necessity for the consideration of multiple CP in networks [19, 20, 21, 22]. One hint from the literature is the fact that the mixture of community structures and CP is interchangeably expressed as ‘CP inside communities’ and ‘communities inside CP’ (see Fig. 1.1 of Ref. [17]). In this subsection, we take both viewpoints by applying the LED to the nodes inside each community (the former) and to coarse-grained communities (the latter). For this (literally) divide-and-conquer strategy, we use the Louvain method [46, 47] to systematically obtain community structures.
@@
First, we implement the LED in each community as shown in Fig. 7(a); we treat each community as an individual network by only considering the nodes and edges inside. Again, we take the CB network [37] and detect communities with the Louvain algorithm [46] using the resolution parameter [47] $\gamma=0.5$ in the modularity function [48]. Although the resultant communities can be different for each realization by the algorithm’s stochasticity [49, 50], we obtain $36$ communities in this case. Then, we take each community (except for $8$ communities with the topology of the star graph, where $S_{cp}$ is undefined because $D_{cc}$ is undefined  [^5] Note that, in a sense, the star graph is the perfect CP and it should strongly positively contribute to $S_{cp}$.) and decompose the nodes inside with the LED, by treating the community and connections inside as a network. In contrast to the core-periphery score $S_{cp}\lesssim 0.1$ at the optimal value $L_{B}^{*}$ for the entire network, the mean value of the core-periphery score averaged over all of the communities is $\langle S_{cp}\rangle_{\mathrm{comm}}=0.32(3)$ (the standard deviation in the parenthesis). Therefore, one can see that the community is an appropriate scale to identify a single CP, by lower structural diversity inside. As an illustrative example with a particularly large $S_{cp}$ value, we take a community with the maximum $S_{cp}(L_{B}^{*}=2)=0.882$. Figure 7(b) shows the connection densities and core-periphery score according to the boundary level, as in the entire network in Fig. 6. To visually inspect the actual organization, we plot the community’s adjacency matrix sorted by the nodes’ hierarchy level in Fig. 7(c). In this case, the optimal value $L_{B}^{*}=2$ happens to be the final decomposition stage, which designates the $3$-clique (marked by the red square) as the core. The well-defined single CP inside a community is observed across most communities, and there are extreme CP cases where a single node forms the core (the star-graph structure) as discussed before.
@@
To check the ‘communities inside CP’ side, we take the coarse-grained point of view and treat each community detected by the Louvain algorithm [46, 47] as a new node or a “supernode” for distinction. The supernode network is composed of supernodes and the edges connecting them if there is any connection between the members of each community, as illustrated in Fig. 8(a) compared to Fig. 7(a); multi-edges and self-loops are ignored for simplicity. Then, on this coarse-grained network [52, 53], we apply the same LED process to detect the CP. Figure 8(b) shows the result for the supernode network from the CB network, again in the case of $\gamma=0.5$. The result indicates $L_{B}^{*}\geq 4$ in this case, which assigns the $7$-clique in the super-adjacency matrix depicted in Fig. 8(c). The super-adjacency matrix looks similar to the ideal CP form indeed; the core nodes are fully connected and most periphery nodes are connected to the core nodes. As in the inside-community version of LED, the structural simplification performed by coarse-graining with Louvain communities enables us to find a simple CP. There are communities or supernodes playing a core role in the super-network composed of supernodes. When using the Louvain method, the number of communities varies depending on the resolution parameter $\gamma$, and the nodes comprising each community change for each run even for a single $\gamma$ value due to its stochasticity. However, except for extreme cases where the resolution is very close to 00 or significantly greater than $1$, results are qualitatively similar to the result obtained in this section.
@@
Finally, we would like to remark on the utility of the hub centrality in this context. The degree as the number of neighbors measures a node’s global significance within the entire network, which is certainly a useful and the most widely used one. However, depending on the purpose of the analysis, comparing the (global) degree values across different focused groups can make deciphering hidden organizational principles even harder, especially in the case of quite heterogeneous degree distributions characterizing real networks [1, 2, 3]. The hub centrality [23, 7] is a tailor-made measure to quantify each node’s relative importance within each focused group composed of the node itself and its neighbors, and it can help us overcome such limitations by standardizing the status of local hubs (recall Fig. 5). Utilizing the hub centrality, the LED assigns standardized levels to local units and identifies the CP based on these levels, and thus, it is versatile across different scales, such as a network, within a community, or a supernode network.
@@ IV Summary and outlook
Naturally, the decomposition yields the core-periphery separation and we have introduced a principled way to pinpoint the boundary by the core-periphery score accounting the similarity to the ideal core-periphery structure with respect to the background edge density; in particular, we would like to emphasize again that our local-edge decomposition shines when it is combined with community division, by preparing the groups with appropriate sizes to handle with the decomposition in advance. We believe that the combination leads to the most natural way to extract so-called multiple core-periphery structures [19, 20, 21, 22], and the extension to the coarse-grained communities as supernodes clearly demonstrates the intermingled structure of core-periphery and communities [17]. From the result, we have clearly shown the merit of embracing local hubs in structurally dissecting networks, as in the previously reported effect on dynamical properties [23, 7]. Most of all, we hope that this type of perspective regarding locally important substructures, such as the concept of hidden dependency between nodes from it [54], gets more attention from the network science community.
@@ Acknowledgments
This work was supported by the National Research Foundation (NRF) of Korea under Grant Nos. NRF-2021R1C1C1004132 (S.H.L.) and NRF-2022R1A4A1030660 (W.J. and S.H.L.). The authors thank Ludvig Lizana for the discussion regarding the concept of using supernodes from network communities in a coarse-grained level.

### v2

@@ III.2 Core-periphery structure of communities and the supernode network
Figure 8: (a) An example of a network with a community structure marked by different background colors. (b) Inside an arbitrary community of the CB network, we plot the connection densities of core-core, core-periphery, and periphery-periphery, along with the core-periphery score, as a function of the boundary level $L_{B}$. The same colors and shapes are used for the densities and the score as in Fig. 7. The core-periphery score $S_{cp}$ reaches its maximum at $L_{B}=9$. (c) The adjacency matrix represents the connections (yellow in the black background) inside the community in panel (b), where nodes are sorted by their LED levels (higher levels toward the left bottom). The nodes with the lowest level are excluded for visibility, and the nodes inside the red square are the optimal set of core nodes of this community.
@@
Figure 9: (a) The supernode network is composed of the nodes corresponding to the communities in the network shown in Fig. 8(a). (b) The connection densities of core-core, core-periphery, and periphery-periphery, along with the core-periphery score, as a function of the boundary level $L_{B}$ in the supernode network from the CB network. The same colors and shapes are used for the densities and the score as in Figs. 7 and 8. The core-periphery score $S_{cp}$ reaches its maximum at $L_{B}=4$. (c) The adjacency matrix represents the connections (green in the black background) between the supernodes in panel (b), where nodes are sorted by their LED levels (higher levels toward the left bottom). The supernodes inside the red square are the optimal core nodes of this community.
@@
In the previous subsection, we have found a single core of the network, but as introduced in the first part of Sec. III, more recent studies on the CP emphasize the necessity for the consideration of multiple CP in networks [19, 20, 21, 22]. One hint from the literature is the fact that the mixture of community structures and CP is interchangeably expressed as ‘CP inside communities’ and ‘communities inside CP’ (see Fig. 1.1 of Ref. [17]). In this subsection, we take both viewpoints by applying the LED to the nodes inside each community (the former) and to coarse-grained communities (the latter). For this (literally) divide-and-conquer strategy, we use the Louvain method [53, 54] to systematically obtain community structures. With our local-hub-based LED scheme applied to each community, we expect to identify core backbone nodes and peripheral shell nodes, which would be closely related to individual nodes’ co-membership consistency with other nodes from stochastic community-detection algorithms (consistent core nodes versus inconsistent peripheral nodes) [55, 56].
@@
First, we implement the LED in each community as shown in Fig. 8(a); we treat each community as an individual network by only considering the nodes and edges inside. Again, we take the CB network [37] and detect communities with the Louvain algorithm [53] using the resolution parameter [54] $\gamma=0.5$ in the modularity function [57]. Although the resultant communities can be different for each realization by the algorithm’s stochasticity [55, 56], we obtain $36$ communities in this case. Then, we take each community (except for $8$ communities with the topology of the star graph, where $S_{cp}$ is undefined because $D_{cc}$ is undefined  [^5] Note that, in a sense, the star graph is the perfect CP and it should strongly positively contribute to $S_{cp}$.) and decompose the nodes inside with the LED, by treating the community and connections inside as a network. Since communities have low structural diversity inside, the distinction between the core and periphery is readily apparent, making it a suitable scale for identifying a single CP. As an illustrative example with a particularly large $S_{cp}$ value, we take a community containing $575$ nodes with the maximum $S_{cp}(L_{B}^{*}=9)=0.378$. Figure 8(b) shows the connection densities and core-periphery score according to the boundary level, as in the entire network in Fig. 7. To visually inspect the actual organization, we plot the community’s adjacency matrix sorted by the nodes’ hierarchy level in Fig. 8(c). In this case, the optimal value $L_{B}^{*}=9$ corresponds to an intermediate decomposition stage. The well-defined single CP inside a community is observed across most communities, and there are extreme CP cases where a single node forms the core (the star-graph structure) as discussed before.
@@
To check the ‘communities inside CP’ side, we take the coarse-grained point of view and treat each community detected by the Louvain algorithm [53, 54] as a new node or a “supernode” for distinction. The supernode network is composed of supernodes and the edges connecting them if there is any connection between the members of each community, as illustrated in Fig. 9(a) compared to Fig. 8(a); multi-edges and self-loops are ignored for simplicity. Then, on this coarse-grained network [59, 60], we apply the same LED process to detect the CP. Figure 9(b) shows the result for the supernode network from the CB network, again in the case of $\gamma=0.5$. The result indicates $L_{B}^{*}=4$ in this case, which assigns the $7$-clique in the super-adjacency matrix depicted in Fig. 9(c). The fact that $L_{B}^{*}$ corresponds to the final stage of decomposition may imply the lack of a characteristic scale or the lack of detailed resolution for this small-sized supernode network composed of $36$ supernodes; the latter hypothesis is supported by the same phenomenon happening for the LED inside a smaller community than the one depicted in Fig. 8. The super-adjacency matrix looks similar to the ideal CP form indeed; the core nodes are fully connected and most periphery nodes are connected to the core nodes. As in the inside-community version of LED, the structural simplification performed by coarse-graining with Louvain communities enables us to find a simple CP. There are communities or supernodes playing a core role in the super-network composed of supernodes. When using the Louvain method, the number of communities varies depending on the resolution parameter $\gamma$, and the nodes comprising each community change for each run even for a single $\gamma$ value due to its stochasticity. However, except for extreme cases where the resolution is very close to 00 or significantly greater than $1$, results are qualitatively similar to the result obtained in this section.
@@
In addition, we would like to remark on the utility of the hub centrality in this context. The degree as the number of neighbors measures a node’s global significance within the entire network, which is certainly a useful and the most widely used one. However, depending on the purpose of the analysis, comparing the (global) degree values across different focused groups can make deciphering hidden organizational principles even harder, especially in the case of quite heterogeneous degree distributions characterizing real networks [1, 2, 3]. The hub centrality [23, 7] is a tailor-made measure to quantify each node’s relative importance within each focused group composed of the node itself and its neighbors, and it can help us overcome such limitations by standardizing the status of local hubs (recall Fig. 6). Utilizing the hub centrality, the LED assigns standardized levels to local units and identifies the CP based on these levels, and thus, it is versatile across different scales, such as a network, within a community, or a supernode network.
III.3 LED versus kCD for a synthetic network with a prescribed CP structure
Figure 10: (a) The number of core nodes and that of all nodes in the remained network with $L\geq L_{B}$ as a function of the boundary level $L_{B}$, from a single realization of the SBM structure in Sec. III.3. The squares and circles represent the results for the LED and kCD, respectively. The filled and open symbols indicate the core and all nodes, respectively. (b) The core-periphery score as a function of the boundary level. The filled squares and circles represent the cases of LED and kCD, respectively.
In this final subsection, we demonstrate the LED’s merit even for a simple structure with gradual coreness. To show a clear advantage of the LED over kCD in this case, we generate a synthetic network using the stochastic block model (SBM) [61]. The synthetic network is composed of three parts denoted by $I_{1}$, $I_{2}$, and $I_{3}$, which are composed of $20$, $40$, and $100$ nodes, respectively. The connection density between different parts is given by $D_{I_{1}I_{1}}=0.7$, $D_{I_{2}I_{2}}=0.3$, $D_{I_{3}I_{3}}=0.1$, $D_{I_{1}I_{2}}=0.5$, $D_{I_{2}I_{3}}=0.2$, and $D_{I_{1}I_{3}}=0$; the nodes inside each group and between groups are connected uniformly at random with the prescribed connection density. The structure represents the three-level coreness, where the primary core part $I_{1}$ do not have any connection to the periphery $I_{3}$ and the secondary core $I_{2}$ plays the role of bridge between the primary core and the periphery. As shown in Fig. 10(a), in the case of kCD-based decomposition, the number of nodes with $L\geq L_{B}$ in the network does not notably change when $L_{B}<10$. In other words, even the level of most peripheral nodes typically goes up to higher than $L_{B}=10$, and only the $60$ nodes constituting the primary and secondary cores remain when $L_{B}\geq 16$. It means that the kCD roughly cuts the structure into two levels: both the primary and the secondary cores as the core and the rest as the periphery. When we apply the LED, in sharp contrast, the number of total nodes with $L\geq L_{B}$ decreases gradually, and in particular, the number of primary and secondary core nodes ($60$ nodes) gradually decreases for $L_{B}>10$, while most decrement in the number of nodes for $L_{B}<10$ is caused by the removal of peripheral nodes ($100$ nodes). Even if we set up the three-level coreness, due to the statistical fluctuation there must be differential levels of local hubs for a single realization of SBM, and the LED successfully captures them as well.
We also measure $S_{cp}$ as a function of $L_{B}$, in the same manner as in Fig. 7. As shown in Fig. 10(b), in the case of the kCD, since most periphery nodes are in high levels ($>10$), $S_{cp}$ remains low across a wide range of lower $L_{B}$ values. At the maximum value of $L_{B}=19$, $S_{cp}$ reaches $0.320$ and the most densely connected $55$ nodes are classified as the core. However, in the case of the LED, $S_{cp}$ starts to increase and reaches its maximum value ($0.453$) at $L_{B}=L_{B}^{*}=16$, and decreases from that point. The number of nodes at $L_{B}^{*}$ is $20$ [Fig. 10(a)], which corresponds to the primary core nodes as expected. In other words, when using the LED, it is possible to identify the most influential nodes even within densely connected nodes. This behavior is consistent with that observed in real networks, shown in Figs. 7 and 8. As $L_{B}$ increases, the proportion of nodes belonging to $I_{1}$ increases among the nodes with $L>L_{B}$, leading to a reduction in the number of connections between nodes with $L\geq L_{B}$ and those with $L<L_{B}$. This results in a lower value of $N_{cp}/N_{p}$ and the presence of the optimal boundary level. Consequently, compared with the kCD, the LED allows the identification of more important nodes and a more detailed analysis even within densely connected structures, such as communities.
@@ IV Summary and outlook
Naturally, the decomposition yields the core-periphery separation and we have introduced a principled way to pinpoint the boundary by the core-periphery score accounting the similarity to the ideal core-periphery structure with respect to the background edge density; in particular, we would like to emphasize again that our local-edge decomposition shines when it is combined with community division, by preparing the groups with appropriate sizes to handle with the decomposition in advance. We believe that the combination leads to the most natural way to extract so-called multiple core-periphery structures [19, 20, 21, 22], and the extension to the coarse-grained communities as supernodes clearly demonstrates the intermingled structure of core-periphery and communities [17]. From the result, we have clearly shown the merit of embracing local hubs in structurally dissecting networks, as in the previously reported effect on dynamical properties [23, 7]. Some related previous works by others include Ref. [38] on community structures, where so-called “whiskers” can be recognized as part of the periphery despite having relatively high degrees because it is based on relative connections. This approach may align in many aspects with the study that utilizes the hub centrality. Most of all, we hope that this type of perspective regarding locally important substructures, such as the concept of hidden dependency between nodes from it [62], gets more attention from the network science community.
@@ Acknowledgments
This work was supported by the National Research Foundation (NRF) of Korea under Grant Nos. NRF-2021R1C1C1004132 (S.H.L.) and NRF-2022R1A4A1030660 (W.J. and S.H.L.). The authors thank Ludvig Lizana for the discussion regarding the concept of using supernodes from network communities in a coarse-grained level, and Heetae Kim for insightful comments including correcting typos.